OpenStudy (anonymous):

Find the equation of the circle passing through the points (4,7) and (8,-1) and having its center on the line x+y=9. thanks

OpenStudy (angela210793):

same logic...

OpenStudy (anonymous):

but I need an answer.

OpenStudy (anonymous):

but I need an answer.

OpenStudy (angela210793):

did u at least try to solve the system i formed?

OpenStudy (anonymous):

pls. just give me the answer... my professor will give a 0.0 grade to me if I haven't answered this 25 items. so far I've already answered 6 items.

OpenStudy (anonymous):

I haven't because I don't know what kind of concept in solving you have given me.

OpenStudy (anonymous):

I think Ishaan has the right idea. Use the formula d^2 = (2r)^2 = 4r^2 = (8-1)^2 + (-1-7)^2 to get r^2. Then use one of the same points in the circle centred at (a,b), (x-a)^2 + (y-b)^2 = r^2 together with 5a-2b = 19 to get (a,b).

OpenStudy (anonymous):

I think using distance formula in solving the equation by the use of the given points is not applicable since the given point is placed on the circumference of the circle... and i think we can only use distance formula in finding the radius by the given point and the center. but i don't know hats next.

OpenStudy (anonymous):

Absolutely, right... Here is a way, get the equation of the line between the 2 points on the circle and find the midpoint. The perpendicular through this midpoint goes through the circle centre, get the equation of that line and solve simultaneously with x+y=9 for the centre. Together with the radius, you are done.